OpenStudy (anonymous):

Can somebody please help me with this quiz? I need somebody to explain the grouping method of factoring, and walk me through a problem when the grouping method would be preferred over other methods.

OpenStudy (anonymous):

@SolomonZelman Can you help me with this please?

OpenStudy (solomonzelman):

grouping? it's not harder than taking 2 numbers out into prime or simpler, numbers and comparing them for example if you had \[\frac{64}{24}\]and you wanted to reduce that, then you would say \[\frac{64}{24}=\frac{2 \times 2 \times 2 \times 2 \times 2 \times 2}{2 \times 2 \times 2 \times 3 }\] by definition, but you can use group factoring like this (doesn't have to be prime numbers when you use it) \[\frac{64}{24}=\frac{8 \times 8}{8 \times 3}~~~~~~8's~~~can cel~~~out.~~~~~->~~~~\frac{3}{8}\]

OpenStudy (solomonzelman):

when you have really huge numbers and you need to reduce then,. then you would rather do grouping.

OpenStudy (anonymous):

Okay, I understand now. Thank You